cone l has a volume of 24 pi cubic inches. If a plane cuts cone l parallel to its base at the midpoint of its altitude, what is volume , in cubic inches , of the small cone formed when the frustum is removed? Express your answer in terms of pi
( a frustum of a cone is the part that is left when the cone is cut by a plane parallel to the base and the part containing the vertex is removed)

Answer by Alan3354(69443)

The small cone's height and radius will be 1/2 of the original.
V = pi*r^2*h/3
The volume is (1/2)^3 of the original, = 1/8
V = 3pi cubic inches